# From Counterfactuals to Trees: Competitive Analysis of Model Extraction Attacks

### v2

@@ 5 Related Works
Tramèr et al. [2016] propose the only functionally equivalent model extraction attack targeting regression or decision trees: PathFinding. It assumes that each query reply contains a unique identifier for the associated leaf. In a nutshell, PathFinding identifies the decision boundaries of each leaf in the target tree by varying the values of each feature. While effective, this method requires a large number of queries, though partial input queries can sometimes mitigate this overhead.
While many recent works have focused on generating counterfactual explanations [Guidotti, 2024], these explanation techniques have also been shown to facilitate privacy attacks [Pawelczyk et al., 2023]. Aïvodji et al. [2020] introduce CF, a model extraction attack that leverages counterfactual explanations. Their approach constructs a labeled dataset by querying both predictions and counterfactuals from the target model, which is then used to train a surrogate. Wang et al. [2022] extend this method with DualCF, which improves fidelity by additionally querying the counterfactuals of the counterfactual explanations. Dissanayake and Dutta [2024] employ polytope theory to show that a sufficient number of optimal counterfactual explanations can approximate convex decision boundaries. They propose a model extraction attack against locally Lipschitz continuous models, with fidelity guarantees dependent on the Lipschitz constants of the target and surrogate models. However, functional equivalence cannot be strictly certified, and as the authors acknowledge, these guarantees do not apply to decision tree-based models, which lack local Lipschitz continuity and convexity.
@@ 6 Conclusions and Discussion
We introduced the first functionally equivalent model extraction attack against decision trees and tree ensembles, leveraging locall optimal counterfactual explanations. In addition to its rigorous functional equivalence guarantee, the proposed method achieves higher fidelity than prior approaches while requiring fewer queries. We also leveraged well-established tools from online discovery to enable a formal analysis of model extraction, drawing an analogy between the two fields. We illustrated the applicability and relevance of this analysis by providing bounds on our attack’s efficiency compared to the best achievable strategy, relying on the notion of competitive ratio. This perspective is essential for formally characterizing and comparing model extraction attacks.
Our study demonstrates that optimal counterfactual explanations can be systematically exploited to reconstruct tree ensembles via query APIs, as they inherently reveal decision boundaries. This raises significant concerns, especially as explainability is increasingly mandated by regulations. In many real-world applications, counterfactual explanations serve as a natural mechanism to meet transparency requirements by providing recourse information. We discuss the broader societal impacts of our work in Appendix D.
@@
Acknowledgements
This research was enabled by support provided by Calcul Québec and the Digital Research Alliance of Canada, as well as funding from the SCALE AI Chair in Data-Driven Supply Chains.
@@ Appendix A Proofs
Applying the Karush-Kuhn-Tucker (KKT) conditions, we find that the maximum occurs when $s^{(n)}_{i}=\frac{n}{m}$ for all $i$. Substituting back, the worst-case complexity becomes $O\left(\left(\frac{n}{m}+1\right)^{m}\right)$. ∎
@@
Let $n>0$, $m\geq 1$, denote $\alpha_{1},...,\alpha_{n}$ the tree split levels (decision thresholds) and for each feature $j=1,...,m$, let $s_{j}$ represent the number of splits in the $j$-th dimension, ordered such that $s_{1}\geq s_{2}\geq\dots\geq s_{m}$. Without loss of generality, assume that the split levels are grouped by dimension. Specifically, splits $\alpha_{1}$ to $\alpha_{s_{1}}$ occur in the first dimension, splits $\alpha_{s_{1}+1}$ to $\alpha_{s_{1}+s_{2}}$ in the second dimension, and so on. Additionally, within each dimension, the split levels are sorted in increasing order, i.e.,
@@
    Proof of Upper Bound:
    In the best-case scenario, where each split level appears exactly once in the decision tree (i.e., there is no redundancy among the tree nodes), any algorithm, including the optimal one, would require at least $12n+$ 1 queries to reconstruct the tree:
    $Q_{opt}^{f}\geq 2n+1$ (2)
    This includes one query for discovering each split and additional queries to verify the leaf nodes.
@@
    $C_{TRA}^{(n,m)}=\text{sup}_{f\in\mathcal{F}}\left(\frac{Q_{TRA}^{f}}{Q_{opt}^{f}}\right)\leq\frac{\sum_{i=1}^{m}s_{i}\prod_{j=1}^{i-1}(s_{j}+1)+\prod_{j=1}^{m}(s_{j}+1)}{2n+1}$ (3)
@@
    where $\epsilon>0$. This adversarial example ensures that within any hyper-rectangle defined by split level boundaries, for $j=1,...,m-1$ if there are splits in both the $j$-th and $j+1$-th dimensions, then there exists a split in the $j$-th dimension that is closer to the center of the hyper-rectangle than any split in the $j+1$-th dimension. As a consequence, TRA will always detect the splits of the $j$-th dimension before those of the $j+1$-th dimension. Therefore, the adversary can design a decision tree with a single branch (as illustrated in the right tree of Figure 6) that begins by splitting on the split levels in reverse (decreasing) order of dimensions (see a 2D example in Figure 6). For this specific example, TRA will require
    $\sum_{i=1}^{m}s_{i}\prod_{j=1}^{i-1}(s_{j}+1)+\prod_{j=1}^{m}(s_{j}+1)$
    queries, whereas the optimal offline algorithm only needs $12n+$ 1 queries.
@@
Figure 6: Adversarial example for TRA (displayed for $m=2$ dimensions). The classes are $c_{1}$,$c_{2}$ and $c_{3}$. For simplicity, we denote $s_{1}=s^{(n)}_{1}$ and $s_{2}=s^{(n)}_{2}$. Here, the instance is dynamically built so that the number of queries required by TRA is $s_{2}(s_{1}+1)+s_{1}+(s_{1}+1)(s_{2}+1)$, whereas the optimal offline algorithm only needs $12n+$ 1, as shown in the right tree figure.
@@
    In this adversarial example, the D&C algorithm requires at least 7 queries to reconstruct the exact decision tree, whereas an optimal algorithm can achieve this with only 5 queries. Therefore, for this adversarial example, no D&C-based algorithm can attain a competitive ratio better than $C_{TRA}^{n,m}=C_{TRA}^{2,2}=\frac{7}{5}$.
@@
    $C^{(n,m)}_{D\&C}\geq\frac{(2n+1)C_{TRA}^{(n,m)}}{2n+1}=C_{TRA}^{(n,m)}$
Appendix B Additional Experimental Results
B.1 Experimental Setup Details
@@
B.2 Configuration of Surrogate-Based Attacks
@@
Knowledge of Target Model Architecture and Hyperparameters.
@@
One first important trend that is consistent across both CF and DualCF, and for both decision trees and random forests, is that knowledge of the hypothesis class of the target model helps fitting a surrogate with high fidelity. Indeed, as can be observed in Figures 8, 9 and 10, the MLP surrogate always under-performs compared to the decision trees or random forests ones. Indeed, fitting a surrogate model of the same type is facilitated by the fact that the shapes of its decision boundaries are the same as the target model, e.g., axis-parallel splits for tree-based models. Interestingly, knowledge of the hyperparameters of the target decision tree or random forests does not seem to help fitting the surrogate. Indeed, in most experiments, the surrogate sharing the same hypothesis class as the target model (i.e., DT or RF) and the surrogate sharing both the hypothesis class and the hyperparameters (i.e., DT+ or RF+) have very close performances. Furthermore, imposing the target model’s hyperparameters to the trained surrogate can even be counterproductive, as can be seen in Figures 8(a) and 8(b) for instance. In such cases, the fact that surrogate learning is more constrained due to the enforced hyperparameters (e.g., maximum depth) seems to slow its convergence towards very high fidelity values. This is consistent with previous findings: for instance, Aïvodji et al. [2020] observed that knowledge of the architecture of a target neural network did not provide a significant advantage to the CF model extraction attack.
Optimality of Counterfactuals.
@@
The results in Figures 8, 9 and 10 suggest that the non-optimal counterfactual explanations returned by DiCE helped fitting the extracted surrogate models better than the optimal ones provided by OCEAN. Indeed, for a fixed query budget and surrogate type, the performances of the extracted model are often better when fitted with DiCE counterfactuals than with OCEAN ones. Although some variations appear, this finding is generally verified for all the studied types of surrogates, for both decision trees and random forests target models, and for both the CF and DualCF extraction attacks. Intuitively, this can be attributed to a greater diversity in the non-optimal counterfactuals, which do not necessarily lie close to a decision boundary, unlike optimal ones. This also highlights a crucial insight: optimal counterfactuals only give an advantage to a model extraction attack if the attack is able to leverage the information it carries as a whole (including both the counterfactual example and its optimality) through a structured approach, as demonstrated by TRA.
B.3 Test Set Fidelity Results
@@
As mentioned in Section 4.1, the results provided in Section 4.2 measure fidelity on a dataset uniformly sampled over the input space, which accurately quantifies how well the extracted models fit the decision boundaries of the target ones over the whole input space. Another approach consists in evaluating fidelity on a test set. In such cases, the results indicate how well the extracted models mimic the target ones for examples drawn from the actual data distribution. We report in Table 2 the results of our extraction attacks against random forests. More precisely, for random forests of varying sizes, we report the average fidelity (measured on the uniformly sampled dataset or on a test set) achieved by all considered methods along with the required number of queries. For DualCF and CF, these values are arbitrarily fixed to allow their surrogates to converge towards (near) perfect fidelity. For our proposed TRA, functional equivalence is achieved using the reported number of queries, hence fidelity on both considered datasets is always $1.0$. For CF and DualCF, we report results for the random forest surrogate using default parameters, for both studied counterfactual oracles. Indeed, we observed in Section B.2 that DiCE counterfactuals led to better anytime performances (in terms of uniform dataset fidelity) than OCEAN ones within the CF model extraction attack, in most experiments. However, this is not always the case, with a few setups in which the difference between both approaches becomes very small or shifts in favor of the runs using OCEAN counterfactuals after sufficiently many iterations. This is the case on the COMPAS dataset (Figure 8(b)), and although the difference remains very subtle, it is also visible on the test set fidelity, illustrating the fact that the two values are often very well aligned.
@@
B.4 Detailed Experimental Results
@@
Figure 12 focuses on functionally equivalent model extraction attacks, and relates the number of queries they require to fully recover the target model to its size (quantified as its number of nodes). The logarithmic scale of the y-axis highlights that TRA usually requires orders of magnitudes less queries than PathFinding to entirely extract a given decision tree, confirming our Result 2 (Section 4.2). Interestingly, this trend is more subtle when reconstructing large trees trained on the datasets with the highest numbers of features (i.e., Adult and SPerformance).
@@
Appendix C Performances of TRA when used with Locally Optimal Counterfactuals
@@
3:  $\mathcal{D}_{\mathcal{E}}\leftarrow\mathcal{D}_{T}\cap\mathcal{E}$gets all the data points that are in the desired input space $\mathcal{E}$}
@@
6:        return  $linesearch(x,x^{\prime})$$x^{\prime}$ is a counterfactual, we refine it via line search towards $x$}
@@
12:        return  $linesearch(x,x^{\prime})$$x^{\prime}$ is a counterfactual, we refine it via line search towards $x$}
@@
15:  return  $None$no counterfactual found}
More precisely, Algorithm 2 searches in the training dataset and then refines the first valid counterfactual it finds via a one-dimensional line search (line 6). If no counterfactual is found, it tries another strategy by sampling a predefined number of points uniformly in $\mathcal{E}$. In our experiments, the maximum number of such sampled points is fixed to $S=1000$. If a valid counterfactual is found, it is refined via line search and returned (line 12). Because the search is stochastic and local, it may fail to return a counterfactual even when one exists, yet it satisfies local optimality whenever a solution is found.
Figures 14 and 15 contrast the performance of TRA to extract decision trees, using OCEAN (globally optimal counterfactuals) versus Algorithm 2 (simpler algorithm, producing locally optimal counterfactuals) as counterfactual oracle. The observed trends confirm show that TRA remains effective as long as the returned counterfactuals are locally, though not necessarily globally, optimal – thereby confirming Result 4 (Section 4.2). Interestingly, we observe in Figure 14 that in the early iterations, locally optimal counterfactuals may lead to better anytime fidelity values – which is likely the case due to their heuristic nature, enhancing diversity. We also observe a pathological case in Figure 14(a), with the Adult dataset. In this particular experiment, when used with Algorithm 2 as counterfactual oracle, TRA converges close to near-perfect – but not exact – fidelity. This is due to the fact that Algorithm 2 fails to find feasible counterfactuals within the desired region, even if one exists. Indeed, even if the counterfactuals provided by this simple oracle satisfy local optimality whenever one is found, there is no guarantee if none is found that none actually exists. Performing a finer uniform sampling (via raising the number of sampled points $S$) could fix this issue. Nevertheless, the performances of TRA with this weak oracle remain very robust, as evidenced through our large set of experiments.
The performance of TRA to extract random forests target models, with either the OCEAN or Algorithm 2 counterfactual oracles, are provided in Table 3 and display the same trends. TRA is able to achieve perfect fidelity for both counterfactual oracles, and there is no significant difference in the required number of queries (which remains very low compared to the size of the reconstructed forests, as discussed in Section 4.2).
@@
Appendix D Broader Impact Statement

### v3

@@ 5 Related Works
Tramèr et al. [2016] propose the only functionally equivalent model extraction attack targeting regression or decision trees: PathFinding. It assumes that each query reply contains a unique identifier for the associated leaf. In a nutshell, PathFinding identifies the decision boundaries of each leaf in the target tree by varying the values of each feature. While effective, this method requires a large number of queries, though partial input queries can sometimes mitigate this overhead. In contrast, TRA does not make strong assumptions regarding the target model’s prediction API, is able to extract any axis-parallel decision boundary model (beyond decision trees), and uses orders of magnitude fewer queries by exploiting counterfactual explanations.
While many recent works have focused on generating counterfactual explanations [Guidotti, 2024], these explanation techniques have also been shown to facilitate privacy attacks [Pawelczyk et al., 2023]. Aïvodji et al. [2020] introduce CF, a model extraction attack that leverages counterfactual explanations. Their approach constructs a labeled dataset by querying both predictions and counterfactuals from the target model, which is then used to train a surrogate. Wang et al. [2022] extend this method with DualCF, which improves fidelity by additionally querying the counterfactuals of the counterfactual explanations. However, neither CF nor DualCF provide fidelity guarantees, and they also do not leverage the structural properties of the target model. Dissanayake and Dutta [2024] employ polytope theory to show that a sufficient number of optimal counterfactual explanations can approximate convex decision boundaries. They propose a model extraction attack against locally Lipschitz continuous models, with fidelity guarantees dependent on the Lipschitz constants of the target and surrogate models. However, functional equivalence cannot be strictly certified, and as the authors acknowledge, these guarantees do not apply to axis-parallel models (such as decision trees), which lack local Lipschitz continuity and convexity. Also note that their approach relies on globally optimal counterfactuals (whereas TRA accommodates locally optimal ones).
@@ 6 Conclusions and Discussion
We introduced the first functionally equivalent model extraction attack against decision trees and tree ensembles, leveraging locally optimal counterfactual explanations. In addition to its rigorous functional equivalence guarantee, the proposed method achieves higher fidelity than prior approaches while requiring fewer queries. We also leveraged well-established tools from online discovery to enable a formal analysis of model extraction, drawing an analogy between the two fields. We illustrated the applicability and relevance of this analysis by providing bounds on our attack’s efficiency compared to the best achievable strategy, relying on the notion of competitive ratio. This perspective is essential for formally characterizing and comparing model extraction attacks.
Our study demonstrates that optimal counterfactual explanations can be systematically exploited to reconstruct tree ensembles via query APIs, as they inherently reveal decision boundaries. This raises significant concerns, especially as explainability is increasingly mandated by regulations. In many real-world applications, counterfactual explanations serve as a natural mechanism to meet transparency requirements by providing recourse information. We discuss the broader societal impacts of our work in Appendix E.
@@
Acknowledgments
This research was enabled by support provided by Calcul Québec and the Digital Research Alliance of Canada, as well as funding from the SCALE AI Chair in Data-Driven Supply Chains. It was also supported by the Fonds de recherche du QuébecNature et technologies (FRQNT) through a Team Research Project (327090).
@@ Appendix A Proofs
Applying the Karush-Kuhn-Tucker (KKT) conditions [Boyd and Vandenberghe, 2004], we find that the maximum occurs when $s^{(n)}_{i}=\frac{n}{m}$ for all $i$. Substituting back, the worst-case complexity becomes $O\left(\left(\frac{n}{m}+1\right)^{m}\right)$. ∎
See 3.6
@@
Let $n>0$, $m\geq 1$, denote $\alpha_{1},...,\alpha_{n}$ the tree split levels (decision thresholds) and for each feature j=1,,mj=1,...,m, let $s_{j}$ represent the number of splits in the $j$-th dimension, ordered such that $s_{1}\geq s_{2}\geq\dots\geq s_{m}$. Without loss of generality, assume that the split levels are grouped by dimension. Specifically, splits $\alpha_{1}$ to $\alpha_{s_{1}}$ occur in the first dimension, splits $\alpha_{s_{1}+1}$ to $\alpha_{s_{1}+s_{2}}$ in the second dimension, and so on. Additionally, within each dimension, the split levels are sorted in increasing order, i.e.,
@@
    Proof of Upper Bound: In the best-case scenario, where each split level appears exactly once in the decision tree (i.e., there is no redundancy among the tree nodes), the omniscient algorithm (a.k.a optimal algorithm) would require at least $n+1$ queries to reconstruct the tree:
    $Q_{opt}^{f}\geq n+1$ (2)
    This includes one query for each leaf to verify its label and certify functional equivalence. Note that this assumes that the omniscient algorithm has correctly guessed the location of each split level and directly queried for counterfactuals over each leaf region.
@@
    $C_{TRA}^{(n,m)}=\text{sup}_{f\in\mathcal{F}}\left(\frac{Q_{TRA}^{f}}{Q_{opt}^{f}}\right)\leq\frac{2\prod_{j=1}^{m}(s_{j}+1)-1}{n+1}$ (3)
@@
    where $\epsilon>0$. This adversarial example ensures that within any hyper-rectangle defined by split level boundaries, for j=1,,m1j=1,...,m-1 if there are splits in both the $j$-th and $j+1$-th dimensions, then there exists a split in the $j$-th dimension that is closer to the center of the hyper-rectangle than any split in the $j+1$-th dimension. As a consequence, TRA will always detect the splits of the $j$-th dimension before those of the $j+1$-th dimension. Therefore, the adversary can design a decision tree with a single branch (as illustrated in the right tree of Figure 7) that begins by splitting on the split levels in reverse (decreasing) order of dimensions (see a 2D example in Figure 7). For this specific example, TRA will require
    $2\prod_{j=1}^{m}(s_{j}+1)-1$
    queries, whereas the optimal offline algorithm only needs $n+1$ queries.
@@
Figure 7: Adversarial example for TRA (displayed for $m=2$ dimensions). The classes are $c_{1}$$c_{2}$ and $c_{3}$. For simplicity, we denote $s_{1}=\smash{s^{(n)}_{1}}$ and $s_{2}=\smash{s^{(n)}_{2}}$. Here, the instance is dynamically built so that the number of queries required by TRA is $s_{2}(s_{1}+1)+s_{1}+(s_{1}+1)(s_{2}+1)$, whereas the optimal offline algorithm only needs $n+1$ queries to check the leaf labels and certify functional equivalence, as shown in the right tree figure.
See 3.7
@@
    In this adversarial example, the D&C algorithm requires at least 7 queries to reconstruct the exact decision tree, whereas an omniscient optimal algorithm can achieve this with only 3 queries, one per leaf. Therefore, for this adversarial example, no D&C-based algorithm can attain a competitive ratio better than $C_{TRA}^{n,m}=C_{TRA}^{2,2}=\frac{7}{3}$.
@@
    $C^{(n,m)}_{D\&C}\geq\frac{(n+1)C_{TRA}^{(n,m)}}{n+1}=C_{TRA}^{(n,m)}.$
Appendix B Details of the Split Procedure (Used by TRA)
Given a region $\mathcal{E}$, a query $x$, and its counterfactual $x^{\prime}$, split partitions $\mathcal{E}$ into disjoint subregions by iterating over features where $x$ and $x^{\prime}$ differ, peeling off the half that contains $x$ and keeping the complementary half (the side toward $x^{\prime}$). The exact pseudo-code of this procedure is provided in Algorithm (2).
Algorithm 2 split(,x,x)($\mathcal{E}$$x^{\prime}$)
0: Region $\mathcal{E}$; vectors $x$, $x^{\prime}$.
0: List $E$ of disjoint subregions whose union is $\mathcal{E}$.
[…]
3:for $(i,v)\in S$ do
4:  if $x_{i}\leq v$ then
5:   $\mathcal{E}_{1}\leftarrow\{\,z\in\mathcal{E}_{0}\;|\;z_{i}\leq v\,\}$
6:  else
7:   $\mathcal{E}_{1}\leftarrow\{\,z\in\mathcal{E}_{0}\;|\;z_{i}>v\,\}$
8:  end if
9:  $E\leftarrow E\cup\{\mathcal{E}_{1}\}$ {peel off the side containing xx}
10:  $\mathcal{E}_{0}\leftarrow\mathcal{E}_{0}\setminus\mathcal{E}_{1}$ {keep the remainder (toward $x^{\prime}$)}
11:end for
[…]
13:return $E$
Tie-handling.
To avoid overlaps at $v$, we use a consistent rule (e.g., $\leq$ on one side and >> on the other). In numerical implementations, we replace the strict inequality ($z_{i}>v$) with a relaxed one $z_{i}\geq v+\varepsilon$ for a small $\varepsilon>0$.
Appendix C Additional Experimental Results
C.1 Experimental Setup Details
@@
C.2 Configuration of Surrogate-Based Attacks
@@
Knowledge of the target model architecture and hyperparameters.
@@
One first important trend that is consistent across both CF and DualCF, and for both decision trees and random forests, is that knowledge of the hypothesis class of the target model helps fitting a surrogate with high fidelity. Indeed, as can be observed in Figures 9, 10 and 11, the MLP surrogate always under-performs compared to the decision trees or random forests ones. Indeed, fitting a surrogate model of the same type is facilitated by the fact that the shapes of its decision boundaries are the same as the target model, e.g., axis-parallel splits for tree-based models. Interestingly, knowledge of the hyperparameters of the target decision tree or random forests does not seem to help fitting the surrogate. Indeed, in most experiments, the surrogate sharing the same hypothesis class as the target model (i.e., DT or RF) and the surrogate sharing both the hypothesis class and the hyperparameters (i.e., DT+ or RF+) have very close performances. Furthermore, imposing the target model’s hyperparameters to the trained surrogate can even be counterproductive, as can be seen in Figures 9(a) and 9(b) for instance. In such cases, the fact that surrogate learning is more constrained due to the enforced hyperparameters (e.g., maximum depth) seems to slow its convergence towards very high fidelity values. This is consistent with previous findings: for instance, Aïvodji et al. [2020] observed that knowledge of the architecture of a target neural network did not provide a significant advantage to the CF model extraction attack.
Optimality of the counterfactuals.
@@
The results in Figures 9, 10 and 11 suggest that the non-optimal counterfactual explanations returned by DiCE helped fitting the extracted surrogate models better than the optimal ones provided by OCEAN. Indeed, for a fixed query budget and surrogate type, the performances of the extracted model are often better when fitted with DiCE counterfactuals than with OCEAN ones. Although some variations appear, this finding is generally verified for all the studied types of surrogates, for both decision trees and random forests target models, and for both the CF and DualCF extraction attacks. Intuitively, this can be attributed to a greater diversity in the non-optimal counterfactuals, which do not necessarily lie close to a decision boundary, unlike optimal ones. This also highlights a crucial insight: optimal counterfactuals only give an advantage to a model extraction attack if the attack is able to leverage the information it carries as a whole (including both the counterfactual example and its optimality) through a structured approach, as demonstrated by TRA.
C.3 Test Set Fidelity Results
@@
As mentioned in Section 4.1, the results provided in Section 4.2 measure fidelity on a dataset uniformly sampled over the input space, which accurately quantifies how well the extracted models fit the decision boundaries of the target ones over the whole input space. Another approach consists in evaluating fidelity on a test set. In such cases, the results indicate how well the extracted models mimic the target ones for examples drawn from the actual data distribution. We report in Table 3 the results of our extraction attacks against random forests. More precisely, for random forests of varying sizes, we report the average fidelity (measured on the uniformly sampled dataset or on a test set) achieved by all considered methods along with the required number of queries. For DualCF and CF, these values are arbitrarily fixed to allow their surrogates to converge towards (near) perfect fidelity. For our proposed TRA, functional equivalence is achieved using the reported number of queries, hence fidelity on both considered datasets is always $1.0$. For CF and DualCF, we report results for the random forest surrogate using default parameters, for both studied counterfactual oracles. Indeed, we observed in Section C.2 that DiCE counterfactuals led to better anytime performances (in terms of uniform dataset fidelity) than OCEAN ones within the CF model extraction attack, in most experiments. However, this is not always the case, with a few setups in which the difference between both approaches becomes very small or shifts in favor of the runs using OCEAN counterfactuals after sufficiently many iterations. This is the case on the COMPAS dataset (Figure 9(b)), and although the difference remains very subtle, it is also visible on the test set fidelity, illustrating the fact that the two values are often very well aligned.
@@
C.4 Detailed Experimental Results
@@
Figure 13 focuses on functionally equivalent model extraction attacks, and relates the number of queries they require to fully recover the target model to its size (quantified as its number of nodes). The logarithmic scale of the y-axis highlights that TRA usually requires orders of magnitudes fewer queries than PathFinding to entirely extract a given decision tree, confirming our Result 2 (Section 4.2). Interestingly, this trend is more subtle when reconstructing large trees trained on the datasets with the highest numbers of features (i.e., Adult and SPerformance).
@@
Appendix D Performances of TRA when used with Locally Optimal Counterfactuals
@@
[…]
@@
6:   return $linesearch(x,x^{\prime})$ $x^{\prime}$ is a counterfactual, we refine it via line search towards xx}
@@
12:   return $linesearch(x,x^{\prime})$ $x^{\prime}$ is a counterfactual, we refine it via line search towards xx}
@@
15:return NoneNone {no counterfactual found}
More precisely, Algorithm 3 searches in the training dataset and then refines the first valid counterfactual it finds via a one‑dimensional line search (line 6). If no counterfactual is found, it tries another strategy by sampling a predefined number of points uniformly in $\mathcal{E}$. In our experiments, the maximum number of such sampled points is fixed to $S=1000$. If a valid counterfactual is found, it is refined via line search and returned (line 12). Because the search is stochastic and local, it may fail to return a counterfactual even when one exists, yet it satisfies local optimality whenever a solution is found.
Figures 15 and 16 contrast the performance of TRA to extract decision trees, using OCEAN (globally optimal counterfactuals) versus Algorithm 3 (simpler algorithm, producing locally optimal counterfactuals) as counterfactual oracle. The observed trends confirm that TRA remains effective as long as the returned counterfactuals are locally, though not necessarily globally, optimal – thereby confirming Result 4 (Section 4.2). Interestingly, we observe in Figure 15 that in the early iterations, locally optimal counterfactuals may lead to better anytime fidelity values – which is likely the case due to their heuristic nature, enhancing diversity. We also observe a pathological case in Figure 15(a), with the Adult dataset. In this particular experiment, when used with Algorithm 3 as counterfactual oracle, TRA converges close to near-perfect – but not exact – fidelity. This is due to the fact that Algorithm 3 fails to find feasible counterfactuals within the desired region, even if one exists. Indeed, even if the counterfactuals provided by this simple oracle satisfy local optimality whenever one is found, there is no guarantee that if none is found that no one actually exists. Performing a finer uniform sampling (via raising the number of sampled points $S$) could fix this issue. Nevertheless, the performances of TRA with this weak oracle remain very robust, as evidenced through our large set of experiments.
The performances of TRA to extract random forests target models, with either the OCEAN or Algorithm 3 counterfactual oracles, are provided in Table 4 and display the same trends. TRA is able to achieve perfect fidelity for both counterfactual oracles, and there is no significant difference in the required number of queries (which remains very low compared to the size of the reconstructed forests, as discussed in Section 4.2).
@@
Appendix E Broader Impact Statement